# Augmenting Molecular Graphs with Geometries via Machine Learning Interatomic Potentials

### v1

@@ Abstract
Accurate molecular property predictions require 3D geometries, which are typically obtained using expensive methods such as density functional theory (DFT). Here, we attempt to obtain molecular geometries by relying solely on machine learning interatomic potential (MLIP) models. To this end, we first curate a large-scale molecular relaxation dataset comprising 3.5 million molecules and 300 million snapshots. Then MLIP foundation models are trained with supervised learning to predict energy and forces given 3D molecular structures. Once trained, we show that the foundation models can be used in different ways to obtain geometries either explicitly or implicitly. First, it can be used to obtain low-energy 3D geometries via geometry optimization, providing relaxed 3D geometries for downstream molecular property predictions. To mitigate potential biases and enhance downstream predictions, we introduce geometry fine-tuning based on the relaxed 3D geometries. Second, the foundation models can be directly fine-tuned for property prediction when ground truth 3D geometries are available. Our results demonstrate that MLIP foundation models trained on relaxation data can provide valuable molecular geometries that benefit property predictions.
@@ 1 Introduction
Molecular property prediction is a critical task in drug discovery, chemistry, and materials science (Zhang et al., 2023; Liyaqat et al., 2024). Many molecular properties are strongly influenced by the stable 3D structure of a molecule, corresponding to its lowest potential energy configuration. For example, as shown in Table 1, in predicting the HOMO-LUMO gap property, GIN (Hu et al., 2021)—which uses only 2D molecular graphs as input—achieves notably worse performance compared to PaiNN (Schütt et al., 2021), which leverages stable 3D geometries. With accurate 3D structures, 3D geometric neural networks (3DGNNs) can significantly improve property prediction accuracy. However, the current standard for obtaining stable molecular structures relies on computationally expensive methods such as density functional theory (DFT) for geometry optimization. Uni-Mol+ (Lu et al., 2023) attempts to bridge this gap by predicting stable 3D geometries during training, allowing only non-stable molecule structures to be used during testing; however, it outperforms GIN but still exhibits a significant performance gap compared to 3DGNNs, highlighting that obtaining useful 3D geometries for property prediction remains a major challenge.
@@
To bridge the gap of obtaining 3D geometries for molecular property prediction, we aim to train a machine learning interatomic potential (MLIP) foundation model to assist geometry relaxation for downstream tasks where only non-stable molecular structures are available during testing. This approach enables the use of downstream 3DGNNs for property prediction using foundation model-relaxed geometries, referred to as Force2Geo in Table 1. Additionally, the foundation model can be directly fine-tuned on downstream tasks when 3D molecular structures are provided. Foundation models have demonstrated remarkable success in computer vision (Bommasani et al., 2021; Liu et al., 2024) and natural language processing (Brown et al., 2020; Touvron et al., 2023), where pre-training yields transferable representations that significantly enhance downstream performance. However, the development of MLIP foundation models for small molecules has been hindered by the lack of large-scale datasets with DFT-level accurate energy and force labels.
In this work, to address the challenge of efficiently obtaining accurate 3D molecular structures, we curate a large-scale molecular relaxation dataset comprising 3.5 million small molecules and 300 million snapshots with energy and force labels, including 105 million snapshots computed using DFT at the B3LYP/6-31G* level of theory. By leveraging this extensive dataset, we can train an MLIP foundation model that can be used in geometry optimization to obtain approximate 3D geometries, referred to as Force2Geo, providing a cost-effective alternative to conventional quantum methods such as DFT. Additionally, we introduce geometry fine-tuning to enhance the downstream predictive accuracy of 3DGNNs using relaxed 3D structures. Furthermore, the MLIP foundation model can be directly fine-tuned for property prediction when ground truth 3D geometries are available, termed Force2Prop, extending its applicability to a range of downstream tasks.
@@ 2 Method
In Section 2.1, we present our curated large-scale relaxation dataset. In Section 2.2, we describe the MLIP foundation model. In Section 2.3, we outline the geometry optimization process using the foundation model. Finally, in Section 2.4, we discuss geometry fine-tuning for molecular property prediction on MLIP foundation model-relaxed structures.
Figure 1: Overview of the MLIP foundation model training pipeline. The model is pre-trained using our curated large-scale relaxation dataset, which includes atomic numbers, forces, positions, and energies for each snapshot. The pre-trained MLIP foundation model can either be fine-tuned for molecular property prediction when stable 3D geometries are available or employed for geometry optimization to obtain 3D geometries for downstream property prediction.
@@ 2.1 Large-Scale DFT Relaxation Dataset
To train an MLIP foundation model, a large-scale relaxation dataset with energy and force labels is essential. However, such a dataset for small molecules is currently unavailable. To address this gap, we curated PubChemQCR (Fu et al., 2025), a new dataset containing DFT-based geometry optimization trajectories for approximately 3.5 million molecules. These molecules are sourced from the PubChem Compound database. The raw trajectory data originate from the PubChemQC database (Nakata and Shimazaki, 2017), and molecular relaxation is performed sequentially using PM3 semi-emperical method, Hartree-Fock, and DFT at the B3LYP/6-31G* level. We extracted atomic numbers, energies, and atomic forces for each snapshot, resulting in 3,471,000 trajectories and 298,751,667 molecular snapshots, of which 105,494,671 snapshots are DFT-calculated. On average, each molecule contains 29 atoms, including 14 heavy atoms. Further dataset details can be found in the PubChemQCR (Fu et al., 2025).
2.2 MLIP Foundation Models for Small Molecules
@@
Machine learning interatomic potentials (MLIPs) are designed to approximate the potential energy surface (PES) of molecular systems, which is traditionally computed using quantum mechanical methods such as density functional theory (DFT). These methods are computationally intensive, motivating the use of MLIPs to learn the PES from DFT-calculated data. The total energy $E$ is predicted based on atomic coordinates $\{\bm{x}_{i}\}_{i=1}^{N}$ and atomic numbers $\{\bm{a}_{i}\}_{i=1}^{N}$, often decomposed as a sum over atom-wise contributions, $E=\sum_{i}E_{i}$. To ensure energy conservation, atomic forces are obtained as the negative gradient of the predicted energy with respect to atomic positions, $\bm{f}_{i}=-\nabla_{\bm{x}_{i}}E$. The pipeline of the training and usage of the MLIP foundation model is shown in Fig. 1. The MLIP foundation model enables efficient geometry optimization to obtain 3D geometries for downstream predictors requiring 3D molecular structures as inputs. Additionally, by capturing the underlying physics of atomic interactions during pre-training, the MLIP foundation model learns informative molecular representations that can be directly fine-tuned for various downstream tasks.
@@
To effectively serve as an MLIP foundation model, a suitable backbone architecture is required to encode molecular information and learn geometric relationships. For molecule representation, we usually represent molecules as graphs $\mathcal{G}=\{V,X,E\}$, where $V\in\mathbb{R}^{n\times d}$ denotes node features, $X\in\mathbb{R}^{n\times 3}$ represents the 3D coordinates of atoms, and $E\in\{0,1\}^{n\times n}$ is the adjacency matrix. In 3D molecular graphs, edges are often constructed using a radius graph, where an edge is formed between two atoms if their Euclidean distance is within a predefined cutoff. Molecules that share the same chemical graph but differ in their 3D coordinates $X$ are referred to as conformers. For backbone models, geometric neural networks are well-suited as they are designed for learning on data with underlying spatial or geometric structures, and are widely used for modeling molecular systems. In this work, we consider existing geometric neural networks as candidate backbones without developing new architectures, as architecture design is outside the scope of this study.
@@
Formally, given a 3D molecular graph $\mathcal{G}=\{V,X,E\}$, where each node has a feature vector $\bm{v}_{i}\in\mathbb{R}^{d}$ and position $\bm{x}_{i}\in\mathbb{R}^{3}$, and each edge has a feature $\bm{e}_{ij}\in\mathbb{R}^{d}$, a general message passing layer for molecular systems at layer $l$ is defined as:
$\bm{m}_{ij}^{(l)}$ $=\phi_{m}\left(\bm{v}_{i}^{(l)},\bm{v}_{j}^{(l)},\bm{x}_{i},\bm{x}_{j},\bm{e}_{ij}\right),$ (1)
$\bm{v}_{i}^{(l+1)}$ $=\phi_{u}\left(\bm{v}_{i}^{(l)},\mathrm{AGG}\left(\left\{\bm{m}_{ij}^{(l)}\mid j\in\mathcal{N}(i)\right\}\right)\right),$ (2)
@@
Since the dataset includes trajectories computed at varying levels of quantum accuracy, we use only the snapshots from the DFT substage to train the MLIP foundation model, leaving the exploration of training with mixed levels of accuracy for future work. Additionally, the DFT substage is the most computationally intensive, taking several hours per molecule, whereas the PM3 and Hartree-Fock stages require only a few minutes. The training objective for the MLIP foundation model includes both energy and force prediction, formulated as:
@@
$\mathcal{L}_{F}$ $=\frac{1}{M}\sum_{j=1}^{M}\|\bm{f}_{j}-\hat{\bm{f}_{j}}\|^{2},$ (6)
where $\lambda_{E}$ and $\lambda_{F}$ are weights to balance the energy and force loss terms, $\hat{e_{i}}$ and $e_{i}$ are the predicted and ground truth energies, and $\hat{\bm{f}_{j}}$ and $\bm{f}_{j}$ are the predicted and ground truth forces for each atom.
@@ 2.3 Geometry Optimization
After training the MLIP foundation model, it can be employed to perform geometry optimization to obtain stable 3D molecular structures. The objective of geometry optimization is to adjust the atomic positions to minimize the potential energy of the system while adhering to predefined convergence criteria.
As shown in Fig. 2, the conventional DFT-based relaxation (Nakata and Shimazaki, 2017) process begins by selecting an exchange-correlation energy functional and a basis set. The electronic structure is then determined iteratively using the self-consistent field (SCF) method. Once the SCF loop converges, the energy of the molecule and the corresponding atomic gradients are calculated. These gradients are then used to update atomic positions through optimization algorithms, such as Newton’s method. This process is repeated until the maximum force falls below the predefined convergence threshold. In the MLIP-based relaxation process, the neural network predicts the forces directly, replacing the DFT calculation, while the remaining procedure, including the optimization loop, remains unchanged.
In this work, we utilize the Broyden–Fletcher–Goldfarb–Shanno (BFGS) (Fletcher, 2000) as the optimization algorithm for geometry optimization. BFGS is a quasi-Newton method that approximates the Hessian matrix to efficiently update atomic coordinates and reduce the overall computational cost. The optimization process iteratively adjusts the atomic coordinates $\{\bm{x}_{i}\}_{i=1}^{N}$ to minimize the potential energy $E(\bm{x})$ predicted by the MLIP foundation model. At each step, the atomic forces $\bm{f}_{i}=-\nabla_{\bm{x}_{i}}E$ are computed to guide the atomic displacements. The maximum force is calculated as:
[…]
@@ 2.4 Geometry Fine-Tuning
Figure 3: Overview of geometry fine-tuning. In the pre-training stage, a property predictor is trained on stable 3D molecules for property prediction. This pre-trained predictor is then fine-tuned on 3D molecular structures relaxed by the MLIP foundation model, with both property prediction and geometry alignment losses.
As discussed in Section 1, when accurate 3D molecular structures are available in the test set, 3DGNNs can predict molecular properties more effectively than Uni-Mol+. The MLIP foundation model can be used to obtain 3D geometries through geometry optimization. However, using foundation model-relaxed conformers may introduce errors and biases due to imperfect force prediction. To address this, we introduce geometry fine-tuning to improve the accuracy of 3DGNN predictions based on these relaxed structures. The whole pipeline of geometry fine-tuning is shown in Fig. 3. Specifically, we first pre-train a downstream predictor on the downstream dataset using ground-truth 3D geometries as input. We then relax the downstream training molecules—starting from unstable conformers—using the pre-trained MLIP foundation model, and fine-tune the downstream predictor on these relaxed structures. This allows the predictor to adapt to the geometric distribution produced by the foundation model. During testing, only the foundation model-relaxed geometries are provided to the downstream predictor.
For geometry fine-tuning, we adopt a multi-task learning framework by introducing geometry alignment as an auxiliary task to support the primary objective of property prediction. This auxiliary task is motivated by two key intuitions: (1) learning to predict deviations from the ground-truth geometry encourages the model to attend to subtle 3D structural cues that are critical for accurate property estimation; and (2) during fine-tuning, the model is exposed to relaxed geometries generated by the foundation model, which inevitably differ from the ground-truth geometries seen during pre-training. This introduces a distribution shift, and the auxiliary geometry alignment task helps bridge this gap by encouraging the model to relate the relaxed geometries back to the original ground-truth domain.
To implement the auxiliary task, we adopt a mixed conformer denoising strategy. During training, half of the input structures are ground-truth conformers with added coordinate noise, and the other half are randomly sampled from the relaxation trajectory produced by the MLIP foundation model. In addition to the main property prediction loss, we include a geometry alignment loss based on cosine similarity, which encourages alignment between the predicted and target atomic displacements. Formally,

### v2

@@ Abstract
Accurate molecular property predictions require 3D geometries, which are typically obtained using expensive methods such as density functional theory (DFT). Here, we attempt to obtain molecular geometries by relying solely on machine learning interatomic potential (MLIP) models. To this end, we first curate a large-scale molecular relaxation dataset comprising 3.5 million molecules and 300 million snapshots. Then MLIP pre-trained models are trained with supervised learning to predict energy and forces given 3D molecular structures. Once trained, we show that the pre-trained models can be used in different ways to obtain geometries either explicitly or implicitly. First, it can be used to obtain approximate low-energy 3D geometries via geometry optimization. While these geometries do not consistently reach DFT-level chemical accuracy or convergence, they can still improve downstream performance compared to non-relaxed structures. To mitigate potential biases and enhance downstream predictions, we introduce geometry fine-tuning based on the relaxed 3D geometries. Second, the pre-trained models can be directly fine-tuned for property prediction when ground truth 3D geometries are available. Our results demonstrate that MLIP pre-trained models trained on relaxation data can learn transferable molecular representations to improve downstream molecular property prediction and can provide practically valuable but approximate molecular geometries that benefit property predictions. Our code is publicly available at: https://github.com/divelab/AIRS/.
@@ 1 Introduction
Molecular property prediction is a critical task in drug discovery, chemistry, and materials science (Zhang et al., 2023; Liyaqat et al., 2024). Many molecular properties are strongly influenced by the stable 3D structure of a molecule, corresponding to its lowest potential energy configuration. For example, as shown in Table˜1, in predicting the HOMO-LUMO gap property, GIN (Hu et al., 2021)—which uses only 2D molecular graphs as input—achieves notably worse performance compared to PaiNN (Schütt et al., 2021), which leverages stable 3D geometries. With accurate 3D structures, 3D geometric neural networks (3DGNNs) can significantly improve property prediction accuracy. However, the current standard for obtaining stable molecular structures relies on computationally expensive methods such as density functional theory (DFT) for geometry optimization. Uni-Mol+ (Lu et al., 2023) attempts to bridge this gap by predicting stable 3D geometries during training, allowing only non-stable molecule structures to be used during testing; however, it outperforms GIN but still exhibits a significant performance gap compared to 3DGNNs, highlighting that obtaining useful 3D geometries for property prediction remains a major challenge.
@@
To bridge the gap of obtaining 3D geometries for molecular property prediction, we aim to train a machine learning interatomic potential (MLIP) pre-trained model to assist geometry relaxation for downstream tasks where only non-stable molecular structures are available during testing. This approach enables the use of downstream 3DGNNs for property prediction using pre-trained model-relaxed geometries, referred to as Force2Geo in Table˜1. We emphasize that Force2Geo produces approximate geometries that may not always converge to the true DFT-optimized structures, and its effectiveness depends on the molecular system and downstream task. Additionally, the pre-trained model can be directly fine-tuned on downstream tasks when 3D molecular structures are provided. pre-trained models have demonstrated remarkable success in computer vision (Bommasani et al., 2021; Liu et al., 2024) and natural language processing (Brown et al., 2020; Touvron et al., 2023), where pre-training yields transferable representations that significantly enhance downstream performance. However, the development of MLIP pre-trained models for small molecules has been hindered by the lack of large-scale datasets with DFT-level accurate energy and force labels.
In this work, to address the challenge of efficiently obtaining accurate 3D molecular structures, we curate a large-scale molecular relaxation dataset comprising 3.5 million small molecules and 300 million snapshots with energy and force labels, including 105 million snapshots computed using DFT at the B3LYP/6-31G* level of theory. By leveraging this extensive dataset, we can train an MLIP pre-trained model that can be used in geometry optimization to obtain computationally efficient but approximate 3D geometries, referred to as Force2Geo, providing a cost-effective alternative to conventional quantum methods such as DFT. The choice of backbone models for pre-training can be found in Section˜4.1. Additionally, we introduce geometry fine-tuning to enhance the downstream predictive accuracy of 3DGNNs using relaxed 3D structures. Furthermore, the MLIP pre-trained model can be directly fine-tuned for property prediction when ground truth 3D geometries are available, termed Force2Prop, extending its applicability to a range of downstream tasks.
Our contributions are threefold:
  • We curate a large-scale molecular relaxation dataset with 3.5M molecules and 300M snapshots, including 105M DFT-level energy and force labels, enabling MLIP model pre-training.
  • We show that the pre-trained MLIP model on our dataset can efficiently produce low-energy 3D geometries via geometry optimization for downstream property prediction, and introduce geometry fine-tuning to further improve 3D GNN performance.
  • We demonstrate that the pre-trained model can be directly fine-tuned for molecular property prediction when ground-truth 3D geometries are available, highlighting the effectiveness of pre-trained MLIPs in supporting diverse downstream tasks.
@@ 2 Method
In Section˜2.1, we present our curated large-scale relaxation dataset. In Section˜2.2, we describe the MLIP pre-trained model. In Section˜2.3, we outline the geometry optimization process using the pre-trained model. Finally, in Section˜2.4, we discuss geometry fine-tuning for molecular property prediction on MLIP pre-trained model-relaxed structures.
Figure 1: Overview of the MLIP pre-trained model training pipeline. The model is pre-trained using our curated large-scale relaxation dataset, which includes atomic numbers, forces, positions, and energies for each snapshot. The pre-trained MLIP model can either be fine-tuned for molecular property prediction when stable 3D geometries are available or employed for geometry optimization to obtain 3D geometries for downstream property prediction.
@@ 2.1 Large-Scale DFT Relaxation Dataset
To train an MLIP pre-trained model, a large-scale relaxation dataset with energy and force labels is essential. However, such a dataset for small molecules is currently unavailable. To address this gap, we curated PubChemQCR (Fu et al., 2025), a new dataset containing DFT-based geometry optimization trajectories for approximately 3.5 million molecules. These molecules are sourced from the PubChem Compound database. The raw trajectory data originate from the PubChemQC database (Nakata & Shimazaki, 2017), and molecular relaxation is performed sequentially using PM3 semi-emperical method, Hartree-Fock, and DFT at the B3LYP/6-31G* level. We extracted atomic numbers, energies, and atomic forces for each snapshot, resulting in 3,471,000 trajectories and 298,751,667 molecular snapshots, of which 105,494,671 snapshots are DFT-calculated. On average, each molecule contains 29 atoms, including 14 heavy atoms. Further dataset details can be found in Appendix˜A and PubChemQCR (Fu et al., 2025).
2.2 MLIP pre-trained Models for Small Molecules
@@
Machine learning interatomic potentials (MLIPs) are designed to approximate the potential energy surface (PES) of molecular systems, which is traditionally computed using quantum mechanical methods such as density functional theory (DFT). These methods are computationally intensive, motivating the use of MLIPs to learn the PES from DFT-calculated data. The total energy $E$ is predicted based on atomic coordinates $\{\bm{x}_{i}\}_{i=1}^{N}$ and atomic numbers $\{\bm{a}_{i}\}_{i=1}^{N}$, often decomposed as a sum over atom-wise contributions, $E=\sum_{i}E_{i}$. To ensure energy conservation, atomic forces are obtained as the negative gradient of the predicted energy with respect to atomic positions, $\bm{f}_{i}=-\nabla_{\bm{x}_{i}}E$. The pipeline of the training and usage of the MLIP pre-trained model is shown in Fig.˜1. The MLIP pre-trained model enables efficient geometry optimization to obtain 3D geometries for downstream predictors requiring 3D molecular structures as inputs. Additionally, by capturing the underlying physics of atomic interactions during pre-training, the MLIP pre-trained model learns informative molecular representations that can be directly fine-tuned for various downstream tasks.
@@
To effectively serve as an MLIP pre-trained model, a suitable backbone architecture is required to encode molecular information and learn geometric relationships. For molecule representation, we usually represent molecules as graphs $\mathcal{G}=\{V,X,A\}$, where $V\in\mathbb{R}^{n\times d}$ denotes node features, $X\in\mathbb{R}^{n\times 3}$ represents the 3D coordinates of atoms, and $A\in\{0,1\}^{n\times n}$ is the adjacency matrix. In 3D molecular graphs, edges are often constructed using a radius graph, where an edge is formed between two atoms if their Euclidean distance is within a predefined cutoff. Molecules that share the same chemical graph but differ in their 3D coordinates $X$ are referred to as conformers. For backbone models, geometric neural networks are well-suited as they are designed for learning on data with underlying spatial or geometric structures, and are widely used for modeling molecular systems. In this work, we consider existing geometric neural networks as candidate backbones without developing new architectures, as architecture design is outside the scope of this study.
@@
Formally, given a 3D molecular graph $\mathcal{G}=\{V,X,A\}$, where each node has a feature vector $\bm{v}_{i}\in\mathbb{R}^{d}$ and position $\bm{x}_{i}\in\mathbb{R}^{3}$, and each edge has a feature $\bm{a}_{ij}\in\mathbb{R}^{d}$, a general message passing layer for molecular systems at layer $l$ is defined as:
$\bm{m}_{ij}^{(l)}$ $=\phi_{m}^{l}\left(\bm{v}_{i}^{(l)},\bm{v}_{j}^{(l)},\bm{x}_{i},\bm{x}_{j},\bm{a}_{ij}\right),$ (1)
$\bm{v}_{i}^{(l+1)}$ $=\phi_{u}^{l}\left(\bm{v}_{i}^{(l)},\mathrm{AGG}\left(\left\{\bm{m}_{ij}^{(l)}\mid j\in\mathcal{N}(i)\right\}\right)\right),$ (2)
@@
Since the dataset includes trajectories computed at varying levels of quantum accuracy, we use only the snapshots from the DFT substage to train the MLIP pre-trained model, leaving the exploration of training with mixed levels of accuracy for future work. Additionally, the DFT substage is the most computationally intensive, taking several hours per molecule, whereas the PM3 and Hartree-Fock stages require only a few minutes. The training objective for the MLIP pre-trained model includes both energy and force prediction, formulated as:
@@
$\mathcal{L}_{F}$ $=\sqrt{\frac{1}{M}\sum_{j=1}^{M}\|\bm{f}_{j}-\hat{\bm{f}_{j}}\|^{2}},$ (6)
where N denotes the number of molecules in a batch, M denotes the number of atoms in a batch, $\lambda_{E}$ and $\lambda_{F}$ are weights to balance the energy and force loss terms, $\hat{e_{i}}$ and $e_{i}$ are the predicted and ground truth energies, and $\hat{\bm{f}_{j}}$ and $\bm{f}_{j}$ are the predicted and ground truth forces for each atom.
@@ 2.3 Geometry Optimization
After training the MLIP pre-trained model, it can be employed to perform geometry optimization to obtain stable 3D molecular structures. The objective of geometry optimization is to adjust the atomic positions to minimize the potential energy of the system while adhering to predefined convergence criteria.
As shown in Fig.˜2, the conventional DFT-based relaxation (Nakata & Shimazaki, 2017) process begins by selecting an exchange-correlation energy functional and a basis set. The electronic structure is then determined iteratively using the self-consistent field (SCF) method. Once the SCF loop converges, the energy of the molecule and the corresponding atomic gradients are calculated. These gradients are then used to update atomic positions through optimization algorithms, such as Newton’s method. This process is repeated until the maximum force falls below the predefined convergence threshold. In the MLIP-based relaxation process, the neural network predicts the forces directly, replacing the DFT calculation, while the remaining procedure, including the optimization loop, remains unchanged.
In this work, we utilize the Broyden–Fletcher–Goldfarb–Shanno (BFGS) (Fletcher, 2000) as the optimization algorithm for geometry optimization. BFGS is a quasi-Newton method that approximates the Hessian matrix to efficiently update atomic coordinates and reduce the overall computational cost. The optimization process iteratively adjusts the atomic coordinates $\{\bm{x}_{i}\}_{i=1}^{N}$ to minimize the potential energy $E(\bm{x})$ predicted by the MLIP pre-trained model. At each step, the atomic forces $\bm{f}_{i}=-\nabla_{\bm{x}_{i}}E$ are computed to guide the atomic displacements. The maximum force is calculated as:
$\text{max\_force}=\max\left(\|\bm{f}_{i}\|\right),$ (7)
where $\bm{f}_{i}\in\mathbb{R}^{3}$ is the force vector for atom $i$ and $\|\cdot\|$ denotes L2 norm. The optimization process terminates when one of the following stopping criteria is met:
  • The maximum force, as defined above, falls below a threshold of $0.05\,\text{eV/Å}$.
@@ 2.4 Geometry Fine-Tuning
Figure 3: Overview of geometry fine-tuning. In the pre-training stage, a property predictor is trained on stable 3D molecules for property prediction. This pre-trained predictor is then fine-tuned on 3D molecular structures relaxed by the MLIP pre-trained model, with both property prediction and geometry alignment losses.
As discussed in Section˜1, when accurate 3D molecular structures are available in the test set, 3DGNNs can predict molecular properties more effectively than Uni-Mol+. The MLIP pre-trained model can be used to obtain 3D geometries through geometry optimization. However, using pre-trained model-relaxed conformers may introduce errors and biases due to imperfect force prediction. To address this, we introduce geometry fine-tuning to improve the accuracy of 3DGNN predictions based on these relaxed structures. The whole pipeline of geometry fine-tuning is shown in Fig.˜3. Specifically, we first pre-train a downstream predictor on the downstream dataset using ground-truth 3D geometries as input. We then relax the downstream training molecules—starting from unstable conformers—using the pre-trained MLIP model, and fine-tune the downstream predictor on these relaxed structures. This allows the predictor to adapt to the geometric distribution produced by the pre-trained model. During testing, only the pre-trained model-relaxed geometries are provided to the downstream predictor.
For geometry fine-tuning, we adopt a multi-task learning framework by introducing geometry alignment as an auxiliary task to support the primary objective of property prediction. This auxiliary task is motivated by two key intuitions: (1) learning to predict deviations from the ground-truth geometry encourages the model to attend to subtle 3D structural cues that are critical for accurate property estimation; and (2) during fine-tuning, the model is exposed to relaxed geometries generated by the pre-trained model, which inevitably differ from the ground-truth geometries seen during pre-training. This introduces a distribution shift, and the auxiliary geometry alignment task helps bridge this gap by encouraging the model to relate the relaxed geometries back to the original ground-truth domain.
To implement the auxiliary task, we adopt a mixed conformer denoising strategy. During training, half of the input structures are ground-truth conformers with added coordinate noise, and the other half are randomly sampled from the relaxation trajectory produced by the MLIP pre-trained model. In addition to the main property prediction loss, we include a geometry alignment loss based on cosine similarity, which encourages alignment between the predicted and target atomic displacements. Formally,